# PsyMem: Fine-grained psychological alignment and Explicit Memory Control for Advanced Role-Playing LLMs

### v1

@@ 5 Experiments
In this section, we explain the dataset, baselines, and metrics design in the experimental setup, and discuss the main results. Additional ablation experiments are discussed in Section 5.3.
@@ 5.1 Experimental Setup
To mitigate potential knowledge bias introduced during the pretraining phase, we construct an evaluation dataset for our main experiments by selecting characters and scenes from novels published after June 2024. This design choice aims to ensure that models have had minimal or no exposure to the selected texts during pretraining, thereby reducing reliance on memorized character-specific knowledge. We first score 100 contemporary novels using large language models (LLMs) and then select 50 representative characters spanning diverse genres. For each character, we collect ten scenes directly sourced from the original text.
@@
In our benchmark evaluation, we tested both open-source and proprietary advanced chatbots, each representing the state-of-the-art within their respective frameworks. For general-purpose models, we compared the proposed PsyMem with several prominent baselines, including GPT-4o(Hurst et al., 2024), GPT-3.5-turbo (OpenAI, 2023),Claude-3.5-sonnet(Anthropic, 2024b), Qwen-Max (Team, 2024), Yi-Large (Young et al., 2024), Deepseek-R1(Guo et al., 2025), DeepSeek-V3 (Liu et al., 2024), Llama3.1-8B-instruct(Grattafiori et al., 2024) and Qwen2.5-7B-Instruct (Yang et al., 2024). These models are considered some of the best in the field, offering a broad range of capabilities across various domains. In the category of Role-Playing-focused baselines, we made comparisons with Hunyuan-Role(Tencent Cloud, 2023), Baichuan-NPC-Turbo(Baichuan AI, 2023), and CharacterGLM-6B (Zhou et al., 2023),Higgs-Llama3-70B(AI, 2024),CoSER-70B(Wang et al., 2025), which are specifically designed to handle character-driven interactions and story-based tasks.
@@
With the proposed PsyMem framework, we train our models based on Llama3.1-8B-Instruct and Qwen2.5-7B-Instruct, and denote the trained models as PsyMem-LLama and PsyMem-Qwen, respectively. All models were trained on four A800 80G GPUs for 3 epochs with a global batch size of 128 and a sequence length of 8,192. LoRA was used with a per-device batch size of 2 and 16 gradient accumulation steps. We adopted a cosine learning rate schedule with a warmup ratio of 0.1, and set the learning rate to 1e-4.
Metrics
We utilize GPT-4o to evaluate 50 established characters across 500 scenarios, each scenario comprising 15 dialogue turns. Evaluations are conducted based on the dimensions described in Section 3, with mean scores calculated accordingly.
@@ 5.2 Main Results
We present the main results in Table 2. We extracted 500 real novel scenes to evaluate all models, assessing them from two perspectives: Character Fidelity and Character-independent Capabilities. Notably, in terms of Memory alignment, Proprietary models demonstrate a more reasonable approach—not merely summarizing but leveraging memory to portray characters more authentically. Although Higgs-LLaMA3-70B and CoSER-70B exhibit relatively stronger performance in memory handling compared to other role-play-specialized models, there remains a noticeable gap between them and top-performing general-purpose models. Within the General Baselines, we observe that Proprietary models still significantly outperform open-source models. GPT-4o and DeepSeek-V3, as outstanding proprietary models, exhibit impressive performance in Character-independent Capabilities. Additionally, Baichuan-NPC-Turbo, as a role-play expertise baseline model, surpasses general baselines in self-consistency, indicating that this metric is particularly important for role-playing agents. Furthermore, it is worth mentioning that DeepSeek-R1 and CoSER-70B excel in Character Fidelity.
Empirical results under the PsyMem framework reveal that, relative to their original counterparts, PsyMem-Qwen and PsyMem-LLaMA achieve character fidelity improvements of 2.07% and 3.50%, respectively. In addition, PsyMem-Qwen demonstrates strong performance in the Social & Leadership and Behavioral Decision dimensions, achieving scores of 80.80% and 78.48%, respectively. PsyMem-LLaMA, on the other hand, obtains a score of 81.93% in the Value dimension, surpassing all other evaluated models in this category. Moreover, PsyMem-Qwen, despite having only 7B parameters, surpassed all measured baselines in Character Fidelity. Additionally, it is worth noting that both models exhibited remarkable improvements in Human-likeness. Overall, the PsyMem framework has demonstrated strong effectiveness in our benchmark tests, enabling precise character control through quantitative character attribute metrics and memory mechanisms.
@@
Table 3: Ablation study on memory integration in the PsyMem framework. Per.: Personality, Val.: Values, SL: Social & Leadership, BD: Behavioral Decision, Mem.: Memory.
Metrics Per. Val. SL* BD*
Cosine 0.98 0.96 0.97 0.97
Table 4: Cosine similarity between human and GPT-4o scores. Per.: Personality, Val.: Values, SL: Social & Leadership, BD: Behavioral Decision.
@@ 5.3 Ablation Study
     + PsyMem (Original PD data) 80.51 70.80 89.49 82.20
     + PsyMem (Role-based PD data) 81.13 85.20 89.93 95.60
@@
     + PsyMem (Original PD data) 81.63 86.60 89.45 93.40
     + PsyMem (Role-based PD data) 82.64 87.60 89.64 96.20
@@
Table 5: Ablation study on Role-based (RB) Pure Dove (PD) dataset in the PsyMem framework.
Figure 4: The impact of data size on model performance using the Pure Dove (PD) dataset.
We further analyzed the effectiveness of memory injection and the incorporation of role-based general SFT data in training.
@@
Memory alignment training. To evaluate the impact of memory alignment training on character fidelity, we constructed a memory testing dataset using GPT-4o, designing 1,500 questions across 50 roles in 500 scenarios based on detailed character profiles. We present our results in Table 3, comparing the performance of PsyMem-LLama and PsyMem-Qwen, with and without memory alignment training. Our findings indicate that incorporating memory data in the second training stage significantly enhances the model’s ability to responding based on retrieved memory, rather than merely summarizing relevant content. This improves both character attribute consistency and long-term memory retention. However, we also observe that memory alignment training has a slight negative impact on character attribute alignment. We hypothesize that this may be due to the presence of a small number of role-specific dialogue examples in the memory data, which could slightly weaken the model’s ability to strictly align with predefined character attributes.
Role-Based General SFT Data We investigate the impact of Role-based General SFT Data on Character-Independent Capabilities and Character Fidelity in Table 5 and Figure 4. Table 5 presents the ablation experiment on the use of Role-based General SFT Data, showing that PsyMem-LLama and PsyMem-Qwen helps mitigate the decline in coherence capabilities caused by catastrophic forgetting of general capabilities. Additionally, it leads to improvements in Character Fidelity, Human-likeness, and Consistency.
In Figure 4, we examine the impact of General SFT Data and Role-based General SFT Data on Character Fidelity for PsyMem-Qwen. The results validate the effectiveness of Role-based General SFT Data. Specifically, we observe that the use of General SFT Data negatively impacts the model’s character-specific abilities, whereas training with Role-based General SFT Data achieves optimal performance with 1,500 samples, after which performance begins to decline. Despite this, it consistently outperforms General SFT Data. We hypothesize that the decline in performance with 3,000 samples of Role-based General SFT Data occurs because the dataset’s distribution becomes overly dominated by GPT-generated content, causing an imbalance in role-specific attributes. A balanced amount of Role-based General SFT Data thus helps improve character-playing abilities without leading to such issues.
@@ 6 Conclusion
In this work, we propose a novel framework PsyMem for controlling character role-playing in large language models (LLMs) by integrating two key aspects: character memory and character attributes, both of which align more closely with modern psychological theories. We validate the proposed PsyMem using a large scale specially designed role-playing dataset based on real novels, addressing the gap between generic and well-known characters in role-playing tasks. Our experiments show that PsyMem-Qwen, trained with our alignment across five dimensions, outperforms all baselines mentioned in this paper, demonstrating strong performance in Character-Independent Capabilities as well.
Furthermore, we explore the benefits of incorporating Role-based General SFT Data for role-playing models. Through this work, we aim to encourage future research on role-playing models that consider psychological frameworks for a more nuanced understanding of character behavior, ultimately advancing the development of more realistic and consistent role-playing LLMs.

### v2

@@ 5 Experiments
In this section, we explain the dataset, baselines, and implementation, and discuss the main results and ablation experiments.
@@ 5.1 Experimental Setup
To mitigate potential knowledge bias introduced during the pretraining phase, we construct an evaluation dataset $\text{D}_{\text{RP}}^{\text{eval}}$ for our main experiments by selecting characters and scenes from novels published after June 2024. This design choice aims to ensure that models have had minimal or no exposure to the selected texts during pretraining, thereby reducing reliance on memorized character-specific knowledge.
We first assess 100 contemporary novels across diverse genres, then select 25 high-quality novels based on character complexity, dialogue richness, expression quality, and narrative perspective. From each selected novel, we identify the two most psychologically complex and representative characters, yielding 50 characters with varied psychological profiles. For each character, we extract ten dialogue-rich scenes directly from the original novels, resulting in 500 evaluation scenarios (50 characters × 10 scenes), which constitutes our evaluation set $\text{D}_{\text{RP}}^{\text{eval}}$. The character attributes are derived from a detailed analysis of the novel using GPT-4o. This comprehensive evaluation framework enables robust assessment of role-playing capabilities across diverse character types and narrative contexts. For the memory evaluation, each scenario includes not only the character’s profile and scene information, but also three key components: a meticulously selected character memory excerpt from the memory text, a memory-dependent question, and its authentic answer serving as the reference answer, allowing us to assess all character fidelity dimensions.
@@
Baselines in our evaluation include both open-source and proprietary advanced chatbots, each representing the state-of-the-art within their respective frameworks. For general-purpose models, we compared the proposed PsyMem with several prominent baselines, including GPT-4o (Hurst et al., 2024), GPT-3.5-turbo (OpenAI, 2023),Claude-3.5-sonnet (Anthropic, 2024b), Qwen-Max (Team, 2024), Yi-Large (Young et al., 2024), Deepseek-R1 (Guo et al., 2025), DeepSeek-V3 (Liu et al., 2024), LLaMA3.1-8B-instruct (Grattafiori et al., 2024) and Qwen2.5-7B-Instruct (Yang et al., 2024). These models are considered some of the best in the field, offering a broad range of capabilities across various domains. In the category of Role-Playing-focused baselines, we made comparisons with Hunyuan-Role(Tencent Cloud, 2023), Baichuan-NPC-Turbo(Baichuan AI, 2023), and CharacterGLM-6B (Zhou et al., 2023),Higgs-LLaMA3-70B(AI, 2024),CoSER-70B(Wang et al., 2025), which are specifically designed to handle character-driven interactions and story-based tasks.
@@
With the proposed PsyMem framework, we fine-tune LLaMA3.1-8B-Instruct and Qwen2.5-7B-Instruct for three epochs, and denote the resulting models as PsyMem-LLaMA and PsyMem-Qwen, respectively. Additional training settings and implementation details are provided in the Appendix B.3. Evaluations are conducted according to Section 4.2.
@@ 5.2 Main Results
We present the main results in Table 2. Within the General Baselines, we observe that proprietary models commonly outperform open-source models. Claude-3.5-sonnet and Yi-Large, as outstanding proprietary models, exhibit impressive performance in Character-independent Capabilities. Additionally, Baichuan-NPC-Turbo, as a role-play expertise baseline model, surpasses general baselines in self-consistency, which is particularly important for role-playing agents. Furthermore, it is worth mentioning that DeepSeek-R1 and CoSER-70B excel in Character Fidelity.
Empirical results under the PsyMem framework reveal that, relative to their original counterparts, PsyMem-LLaMA and PsyMem-Qwen achieve character fidelity improvements of 2.07% and 3.50%, respectively. In addition, PsyMem-Qwen demonstrates strong performance in the Social & Leadership and Behavioral Decision dimensions, achieving scores of 80.80% and 78.48%, respectively. PsyMem-LLaMA, on the other hand, obtains a score of 81.93% in the Value dimension, surpassing all other evaluated models in this category. Moreover, PsyMem-Qwen, despite having only 7B parameters, surpassed all measured baselines in Character Fidelity. Additionally, it is worth noting that both models exhibited remarkable improvements in Human-likeness. Overall, the PsyMem framework has demonstrated strong effectiveness in our benchmark tests, enabling precise character control through quantitative character attribute metrics and memory mechanisms.
@@
Table 3: Ablation studies on DMT. and memory integration within the PsyMem framework.
@@ 5.3 Ablation Study
     + PsyMem (PD data) 80.51 70.80 89.49 82.20
     + PsyMem ($\text{D}_{\text{RP}}^{\text{synth}}$) 81.13 85.20 89.93 95.60
@@
     + PsyMem (PD data) 81.63 86.60 89.45 93.40
     + PsyMem ($\text{D}_{\text{RP}}^{\text{synth}}$) 82.64 87.60 89.64 96.20
@@
Table 4: Ablation study on Synthetic Role-playing style ($\text{D}_{\text{RP}}^{\text{synth}}$) Pure Dove (PD) dataset in PsyMem.
Figure 4: Comparing Original and Synthetic Role-Playing data ($\text{D}_{\text{RP}}^{\text{synth}}$). PD: Performance by Dataset Size.
@@
Memory alignment training. To evaluate the effectiveness of memory alignment training, we compare the performance of PsyMem-LLaMA and PsyMem-Qwen with and without memory integration during the second training stage. The results are presented in Table 3. Our findings indicate that incorporating memory data in the second training stage significantly enhances the model’s ability to responding based on retrieved memory, rather than merely summarizing relevant content. This improves both character attribute consistency and long-term memory retention. However, we also observe that memory alignment training has a slight negative impact on character attribute alignment. We hypothesize that this may be due to the inclusion of a limited number of role-specific dialogue examples in the memory data, which could slightly diminish the model’s ability to adhere strictly to predefined character attributes.
Two-stage Training Strategy. To evaluate the effectiveness of the DMT strategy, which serves as the core of our two-stage training strategy approach, we conducted ablation studies comparing the full PsyMem framework (with DMT) against a single-stage fine-tuning baseline, in which all datasets and loss functions are jointly optimized. As shown in Table 3, DMT consistently improves performance across both backbone models (Qwen and LLaMA). For PsyMem-Qwen, DMT leads to an overall improvement of 0.67% (82.64% vs. 81.97%), while for PsyMem-LLaMA, the improvement is 0.49% points (81.13% vs. 80.64%), demonstrating the effectiveness of the two-stage approach in enhancing character fidelity.
Synthetic Role-playing Data. We investigate the impact of Synthetic Role-playing style SFT Data $\text{D}_{\text{RP}}^{\text{synth}}$ on model performance through ablation studies presented in Table 4 and Figure 4.
Table 4 compares model performance using $\text{D}_{\text{RP}}^{\text{synth}}$ versus original general SFT data. The results demonstrate that incorporating $\text{D}_{\text{RP}}^{\text{synth}}$ significantly improves Character Fidelity, Human-likeness, and Consistency for both PsyMem-LLaMA and PsyMem-Qwen, while effectively maintaining coherence capabilities.
Figure 4 further examines the effect of data volume, comparing General SFT Data with $\text{D}_{\text{RP}}^{\text{synth}}$ on role-playing performance. The results validate the superiority of our synthetic approach: while General SFT Data consistently degrades character-specific abilities, Synthetic Role-playing style data achieves optimal performance at 1,500 samples before declining at larger volumes. We attribute this performance drop at 3,000 samples to dataset distribution imbalance, where excessive synthetic content may dilute role-specific attribute diversity.
Figure 5: Ablation study of each dimensions. Per.: Personality, Val.: Values, SL: Social & Leadership, BD: Behavioral Decision, Mem.: Memory.
Character Attribute Dimension. We assess the contribution of each character attribute by conducting an ablation study on PsyMem-Qwen, where each of the four supervision signals—Personality, Values, Social & Leadership, and Behavioral Decision—is individually removed (see Figure 5). The results reveal that removing any single dimension markedly drops performance on its corresponding evaluation metric. This indicates each dimension is critical to the model’s fidelity for that trait, underscoring the necessity of the proposed comprehensive attribute set. We have placed more ablation experiments on dimensions in Appendix D.
@@ 6 Conclusion
This paper presented PsyMem, a novel framework for LLM-based role-playing that synergizes character memory with fine-grained psychological attributes, aligning with contemporary psychological theories. Evaluations conducted on a large-scale novel-based dataset demonstrate PsyMem’s effectiveness in role-playing. The resulting model, PsyMem-Qwen, even with only 7B size parameters, outperforms all baselines in character fidelity while also demonstrating strong character-independent capabilities. Additionally, we highlighted the advantages of integrating Synthetic Role-playing style General SFT Data. We believe this work paves the way for future research into psychologically-informed role-playing models, fostering more nuanced, realistic, and consistent character portrayals and thereby expanding the utility of LLMs in applications such as trustworthy social simulation, enhanced human-computer interaction, and AI counseling.
Appendix
Appendix A Source Dataset
Our dataset was sourced from public English-language novel recommendations on platforms including The StoryGraph (The StoryGraph, 2024), Goodreads (Goodreads, 2024) and BookLikes (BookLikes, 2024). An initial pool of 1,710 novels was compiled and subsequently filtered using a Large Language Model (LLM), resulting in a final selection of 539 novels that satisfy the following criteria:
  • Fluent, reader-friendly novel.
  • Third-person narration.
  • Clear formatting and well-structured layout.
  • Balanced genre coverage to ensure diversity.
Appendix B Implementation Details
B.1 Memory Module
We integrate Memory Module into the PsyMem framework to support contextual memory retrieval during role-playing. Specifically, we employed Nano-GraphRAG (Ye, 2024), an open-source lightweight library that implements a simplified version of GraphRAG (Edge et al., 2024). Since our test set consists of recently published novels, PsyMem relies on GraphRAG during inference to enrich the contextual content of the dialogue. We used a paragraph length of 8192 tokens for the memory text in local mode. Other configurations followed the default settings of Nano-GraphRAG. Specifically, in each dialogue turn, GraphRAG retrieves relevant entity information mentioned in the prior context to search for memory that aligns with the role the model is playing, and then combines this with character attributes to support inference.
B.2 Training Details
We applied the PsyMem method to both Qwen2.5-7B-Instruct and LLaMA3.1-8B-Instruct models. During training, we incorporated data augmentation techniques to enhance the models’ ability to understand and generate content in role-playing tasks. Specifically, for each training instance, there was a 50% probability of shuffling or masking the role attributes to expand the dataset.
B.3 Hyperparameters
All models were trained on four A800 80G GPUs for 3 epochs with a global batch size of 128 and a sequence length of 8,196 (increased to 16,384 during the second training stage). LoRA was used with a per-device batch size of 2 and 16 gradient accumulation steps. We adopted a cosine learning rate schedule with a warmup ratio of 0.1, and set the learning rate to 1e-4.
Appendix C Additional Evaluation
Model Character Fidelity
Per. Val. SL BD Mem. Avg.
PsyMem-LLaMA (Rewrite) 80.14 81.47 78.62 74.35 89.80 80.88
PsyMem-LLaMA (Random) 80.33 81.58 78.37 74.56 90.20 81.00
PsyMem-LLaMA (Original) 80.47 81.93 78.57 74.47 90.20 81.13
PsyMem-Qwen(Rewrite) 80.70 81.36 80.20 78.42 91.00 82.34
PsyMem-Qwen(Random) 80.45 81.63 81.02 78.39 92.00 82.70
PsyMem-Qwen(Original) 80.40 81.74 80.80 78.48 91.80 82.64
Table 5: Evaluation of the model’s generalization to input with freer phrasing (Rewrite), shuffled dimension order (Random), and prompts consistent with those used during training (Original).
C.1 More Flexible Inputs
To assess PsyMem’s robustness to varied prompt styles, we tested three evaluation settings: the original structured prompt (‘Original’), prompts with randomly shuffled character dimensions (‘Random’), and character profiles rewritten into natural paragraphs by GPT-4o (‘Rewrite’). Both PsyMem-LLaMA and PsyMem-Qwen were evaluated across five Character Fidelity dimensions.
As shown in Table 5, the ‘Random’ variant achieves comparable performance to the ‘Original’ version, suggesting that the model is robust to prompt order perturbations (likely a result of training-time augmentation). The ‘Rewrite’ variant results in a slight performance decrease (0.25% for LLaMA, 0.30% for Qwen), potentially due to minor information loss during rephrasing. Nonetheless, consistently high Memory scores across all variants demonstrate the robustness and generalization ability of PsyMem under varying input conditions.
C.2 Evaluation strategies
To verify the robustness of our evaluation methodology, we compared two scoring strategies: one strategy where each character’s dialogues across different scenes were first concatenated and then assessed collectively, and another strategy where each character’s dialogue, consisting of 15 turns per scene, was evaluated separately before being aggregated. We applied both strategies to four top-performing models, using three discriminative evaluators—GPT-4o (Hurst et al., 2024), OpenAI-o3 (OpenAI, 2025), and Gemini-2.5-pro (DeepMind, 2025). As shown in Table 6, results across all models and evaluators exhibit strong consistency between the two strategies. Notably, for PsyMem-Qwen under GPT-4o evaluation, the average Character Fidelity score under concatenated dialogue input was 82.87%, only 0.23% points higher than the segment-wise evaluation score of 82.64%, indicating that the choice of evaluation strategy does not significantly impact overall conclusions.
Model GPT-4o o3 Gemini-2.5-pro
GPT-4o 80.86 79.62 76.92
+ Concatenated Eval. 81.00 79.39 76.70
Deepseek-R1 82.47 78.40 75.87
+ Concatenated Eval. 82.26 78.66 75.99
CoSER-70B 82.28 80.23 79.16
+ Concatenated Eval. 81.99 80.40 78.99
PsyMem-Qwen 82.64 80.74 79.70
+ Concatenated Eval. 82.87 80.43 79.39
Table 6: Average Character Fidelity scores (Per., Val., SL, BD, Mem.) of advanced role-playing models evaluated by different discriminative models and evaluation approaches. “Concatenated Eval.” refers to presenting all dialogues across scenarios simultaneously to the discriminative model for evaluation.
Metrics Per. Val. SL BD
Cosine (H $\leftrightarrow$ L) 0.96 0.95 0.96 0.94
$r_{s}$ (H) 0.74 0.79 0.75 0.77
$\rho$ (H) 0.72 0.71 0.71 0.70
Mean (Std) (H) 0.59 $\pm$ 0.47 0.46 $\pm$ 0.42 0.42 $\pm$ 0.39 0.54 $\pm$ 0.41
$\alpha$ (H) 0.77 0.80 0.74 0.80
Table 7: Cosine similarity, Pearson correlation coefficient ($\rho$), Spearman’s correlation coefficient ($r_{s}$), and Mean (Std) are used to assess the agreement and score distribution between human raters and GPT-4o across different evaluation dimensions. For reference, Krippendorff’s Alpha ($\alpha$) is also reported as a measure of inter-rater reliability among human annotators.
C.3 Human Evaluation
To validate our LLM-based evaluation framework, we conducted a large-scale human study involving 40 volunteers who assessed 25 dialogue scenarios (50 characters). The process was highly time-intensive, requiring over 10 hours per rater. For a robust analysis, we employed several complementary metrics. We used Krippendorff’s Alpha ($\alpha$) to measure inter-rater reliability among human annotators, thereby establishing the reliability of our evaluation baseline. Concurrently, we assessed the agreement between GPT-4o’s scores and the mean human scores using Cosine Similarity, Pearson’s correlation coefficient ($\rho$), and Spearman’s rank correlation coefficient ($r_{s}$).
The results, summarized in Table 7, validate our approach. We first confirmed a reliable human baseline, with Krippendorff’s Alpha ($\alpha$) values of 0.74–0.80 indicating substantial inter-rater reliability (Krippendorff, 2018; Marzi et al., 2024). Against this baseline, GPT-4o’s scores showed strong agreement with human ratings. Pearson’s correlation ($\rho$) was consistently strong (0.70–0.72), and Spearman’s rank correlation ($r_{s}$) was similarly high (0.74–0.79) (Fleiss, 1971; Hauke and Kossowski, 2011). These results collectively demonstrate that GPT-4o serves as a valid proxy for human evaluation in our task.
Appendix D Evaluation-stage Input Ablation
Prompt Set Per. Val. SL BD
All 80.40 81.74 80.80 78.48
Per. 78.35 73.33 74.72 73.67
Val. 74.74 79.66 74.35 74.39
SL* 68.12 67.86 79.64 69.35
BD* 69.38 68.77 70.14 77.76
Base 66.01 65.65 68.05 67.19
Table 8: Performance of PsyMem-Qwen with different prompt combinations across four evaluation dimensions. “All” indicates inclusion of all prompts for the four dimensions; “Base” indicates exclusion of these prompts. Bold: Best; Underlined: Second best.
We further investigated the role of character attributes during inference by varying the input format at evaluation time (as shown Table 8). Specifically, we compared three input conditions: (1) a minimal setting with no character dimensions, (2) inputs with only one active dimension, and (3) full inputs containing all four dimensions.
The model’s performance deteriorated significantly when all dimensions were removed—for example, average scores dropped to as low as 66.01 on the Personality axis and 65.65 on Values. In contrast, even a single active dimension led to improved performance on the corresponding evaluation axis; using only the Personality prompt recovered the Personality score to 78.35, and similarly, Values alone yielded 79.66 on its dimension. Additionally, we observed that providing Personality and Values information boosted performance on Social & Leadership and Behavioral Decision dimensions (74.72 and 73.67 under Personality-only input), while the reverse was not true. This may suggest a hierarchical relationship among these attributes, where latent psychological attributes act as foundational traits that inform and guide performance in the more applied, explicit behavioral patterns.